### 10 Mar 2026 v2

# 0Year 0Year 0Year

Abstract
Motivation: Phenotype concept recognition (CR) is a fundamental task in biomedical text mining. However, existing methods either require ontology-specific training making them struggle to generalize across diverse text styles and evolving biomedical terminology, or depend on general-purpose large language models (LLMs) that lack necessary domain knowledge.
Results: To address these limitations, we propose AutoPCR, a prompt-based phenotype CR method designed to automatically generalize to new ontologies and unseen data without ontology-specific training. To further boost performance, we also introduce an optional self-supervised training strategy. Experiments show that AutoPCR achieves the best average and most robust performance across datasets. Further ablation and transfer studies demonstrate its inductive capability and generalizability to new ontologies.
issue: x
1 Introduction
Biomedical text mining plays a key role in unlocking clinical and scientific knowledge from unstructured data sources such as clinical notes and research articles. A fundamental step in this process is ontology-based concept recognition (CR), which aims to identify textual mentions of concepts defined in a given ontology from input text in a zero-shot setting. An ontology is a formal, structured representation of domain-specific knowledge curated by experts, consisting of standardized concepts with associated names, definitions, synonyms, and hierarchical relationships. Phenotype CR, a specific instance of ontology-based CR using the Human Phenotype Ontology (HPO) (Talapova et al., 2023), has become a central research focus. This is largely due to the availability of richly annotated datasets (Weissenbacher et al., 2023; Lobo et al., 2017; Anazi et al., 2017), yet the task remains challenging because of the specialized and rapidly evolving nature of biomedical terminology and ontologies. Phenotype CR plays a critical role in downstream biomedical applications. For example, genetic disease diagnostics require accurate identification of phenotype concepts in clinical notes (Labbé et al., 2023), and biomedical knowledge graph construction relies on robust CR from scientific literature to support integrative data analysis and knowledge discovery (Huang et al., 2024).
Traditionally, CR can be divided into two stages: (1) entity extraction, which identifies text spans to be linked to ontology concepts; and (2) entity linking, which maps these spans to semantically corresponding ontology concepts. Early CR methods are primarily dictionary-based, which rely on lookup tables and string-matching techniques to identify concept mentions. While these approaches offer high precision, they suffer from low recall due to limited vocabulary coverage and inability to handle linguistic variations (Jonquet et al., 2009; Taboada et al., 2014). In recent years, researchers have increasingly turned to neural methods that fine-tune domain-specific pre-trained language models (PLMs), such as BioBERT (Lee et al., 2020), on the rich synonym sets provided by the ontology. By distilling knowledge from synonymous relationships, these models effectively mitigate the challenges of lexical heterogeneity. However, as they are trained against a fixed ontology, re-training is required to recognize new concepts, limiting their usability for frequently updated ontologies like HPO as well as generalizability to unseen ontologies. In contrast, large language models (LLMs), such as GPT-4 (Achiam et al., 2023), have demonstrated strong zero-shot learning capabilities, offering new possibilities for phenotype CR. Recent studies have shown that LLMs can effectively extract clinical information, where phenotype concepts abound, without domain-specific fine-tuning (Meoni et al., 2023). Despite their potential, LLMs face challenges related to factual consistency and reliability in knowledge-intensive tasks (Reese et al., 2024). To address these challenges, retrieval-augmented generation (Lewis et al., 2020, RAG) has emerged as an effective technique to improve LLM’s CR performance by incorporating relevant information retrieved from the ontology through semantic similarity-based retrieval mechanisms. However, existing RAG-based methods often rely on general-purpose components that fail to capture the nuanced semantics of highly specialized and dynamic biological concepts.
In this study, we propose AutoPCR, an automated phenotype CR framework based on prompting, which consists of three sequential steps. First, to ensure that extracted entities are biologically meaningful with high coverage, AutoPCR employs a unified entity extraction approach including BioNER and syntax-based strategies. In the second step, AutoPCR retrieves candidate concepts using a domain-specific semantic similarity model. Although phenotype CR has been extensively studied, prior work has largely overlooked advances in distantly-supervised biomedical entity linking. AutoPCR adopts one of these models named SapBERT (Liu et al., 2021) to accurately identify semantically relevant concepts. In the final step, AutoPCR performs entity linking by prompting an LLM. Each entity and its associated candidate set are encoded into a structured prompt that includes the synonyms and definition of each candidate concept. This approach enables accurate disambiguation even without ontology-specific fine-tuning. To further enhance its capability in distinguishing similar concepts, we introduce a self-supervised fine-tuning strategy that enables AutoPCR to learn from difficult positive and negative examples drawn by SapBERT. Experiments demonstrate the following advantages of AutoPCR: (1) Superior and robust performance. AutoPCR achieves the best average and most robust performance across various benchmarks, changing the status qua that prompt-based methods are lagging behind dictionary-based and neural methods. (2) Inductive capability. AutoPCR maintains superior performance even without prior exposure to HPO, making it adaptable to the frequent updates of HPO. (3) Generalizability. AutoPCR transfers well to a new ontology without reconfiguration and supports rapid deployment within minutes, offering great potential for broad application.
2 Related work
Phenotype concept recognition
Phenotype CR has grown into multiple methodological paradigms, including dictionary-based, neural, and prompt-based methods.
Dictionary-based methods use lookup tables or string similarity measures to identify entities by exhaustively matching input spans against ontology concepts. Tools such as NCBO (Jonquet et al., 2009), OBO (Taboada et al., 2014), Doc2HPO (Liu et al., 2019), and ClinPhen (Deisseroth et al., 2019) exemplify this strategy. Recently, FastHPOCR (Groza et al., 2024b) utilizes groups of morphologically equivalent words generated by GPT-4 to address lexical variations.
Neural methods leverage deep learning architectures, such as convolutional neural networks (CNNs) and BERT (Devlin et al., 2019), with supervised learning on ontology-derived synonym pairs to improve performance. There are two common entity extraction strategies adopted by neural methods, serving as the foundation for subsequent entity linking. One line of methods, e.g., PhenoTagger (Luo et al., 2021) and PhenoTagger++ (Qi et al., 2024), adopts a syntax-based strategy that exhaustively generates entity spans after removing punctuation and function words. PhenoTagger predicts concepts from extracted entities by fine-tuning a BioBERT. PhenoTagger++ improves PhenoTagger by augmenting training data with LLM-generated synonyms and character edition, as well as introducing graph-based concept embeddings from the ontology structure. The other line of methods use off-the-shelf BioNER tools such as Stanza (Zhang et al., 2021) to identify biomedical segments and then exhaustively extract entities from them. NCR (Arbabi et al., 2019) encodes entities using pre-trained text embeddings followed by a CNN and matches them with learnable embeddings of ontology concepts after hierarchical aggregation. PhenoBERT (Feng et al., 2022) builds upon NCR with multiple CNNs for phenotype subcategories to retrieve candidate concepts and a trainable BioBERT to score each entity-concept pair. Recently, PhenoBCBERT and PhenoGPT (Yang et al., 2024) train Bio+ClinicalBERT (Alsentzer et al., 2019) and GPT-3 (Brown et al., 2020) on manually labeled data to improve rare disease recognition.
LLMs, such as GPT-4, enable phenotype CR without the need of ontology-specific training, allowing for rapid adaptation to evolving biomedical vocabularies. Prompt-based methods, such as Labbé et al. (2023) and Groza et al. (2024a), employs prompt engineering to directly extract biomedical concepts from text. REAL (Shlyk et al., 2024) in a RAG style relies on LLMs to extract biomedical entities and perform entity linking. Candidate concepts are retrieved for each entity using a general-purpose embedding model based on names and definitions, and then passed to an LLM for final linking.
Distantly-supervised biomedical entity linking
Distant supervision has emerged as a practical solution for biomedical entity linking, where manual annotation is costly and often infeasible. Given the Unified Medical Language System (Bodenreider, 2004, UMLS), a comprehensive biomedical ontology with over 4M concepts covering HPO, these methods leverage ontology-derived supervision, enabling large-scale training without labeled corpora. BioSyn (Sung et al., 2020) aligns entities with concept synonyms using a combination of character-level features and learnable embeddings. SapBERT fine-tunes a PubMedBERT (Gu et al., 2021) through contrastive learning on hard concept name-synonym pairs. KrissBERT (Zhang et al., 2022) incorporates concept context from PubMed Canese and Weis (2013) into contrastive learning and employs a cross-attention encoder to re-rank candidate concepts. Its inference strategy retains multiple context embeddings per concept, enabling context-aware matching. Recent work (Sasse et al., 2024) shows that synthetic entities generated by LLMs can further enhance linking performance, especially under distribution shifts, highlighting the complementary role of generative models in distant supervision.
3 Method
Figure 1: Architecture of AutoPCR. It performs concept recognition in three stages: entity extraction using BioNER and syntactic entity extraction (e.g., extracted entities $e_{1}$, $e_{2}$, and $e_{3}$), candidate concept retrieval via SapBERT initialized from PubMedBERT (e.g., retrieved concepts $c_{1}$ and $c_{2}$ for $e_{3}$), and entity linking through prompting an LLM (e.g., linked concept $c_{1}$).
In this section, we present AutoPCR, an automated method for phenotype CR based on prompting. We begin with a formal definition of the CR task and describe how it can be decomposed into sequential subtasks. We then define each subtask rigorously and elaborate on how AutoPCR addresses them through an integrated and modular design. The overall architecture of AutoPCR is shown in Figure 1.
Definition 1 (Concept recognition).
Given an ontology O=(C,I)O{=}(C,I) with concepts $C{=}\{c_{1},\dots,c_{n}\}$ and concept-level information $I$ (e.g., definitions and synonyms), for a piece of input text $T$, the task is to find $f$ such that
$f(O,T)=\{(i,j,c)\mid i,j\in[0,|T|],\;T[i\mathord{:}j]\sim c\in C\}$ (1)
which extracts entities from the input text with start and end offsets that can be mapped to semantically similar concepts in the ontology.
We decompose the CR task $f$ into three sequential subtasks $f_{\text{EL}}\circ f_{\text{CCR}}\circ f_{\text{EE}}$, which are entity extraction $f_{\text{EE}}$, candidate concept retrieval $f_{\text{CCR}}$, and entity linking $f_{\text{EL}}$.
Definition 2 (Entity extraction).
Given an ontology O=(C,I)O{=}(C,I) and a piece of input text $T$, the task is to find $f_{\text{EE}}$ such that
$f_{\text{EE}}(O,T)\supset\{(i,j)\mid i,j\in[0,|T|],\;T[i\mathord{:}j]\sim\exists c\in C\}$ (2)
which extracts entity spans from the text that may correspond to ontology concepts.
We employ a unified entity extraction approach that combines BioNER and syntax-based strategies, making it applicable to both free-form clinical notes and standardized biomedical abstracts. For BioNER, we follow PhenoBERT and process each sentence using Stanza’s BioNER model to extract biomedically relevant segments. To improve coverage, additional segments are generated by splitting on punctuation and conjunctions, and all resulting segments are used as windows for $n$-gram entity extraction ($n{\in}[2,10]$). To better cover free-form text, we complement BioNER with a syntactic entity extraction component, which identifies structurally coherent phrases through noun phrase extraction, coordinated phrase decomposition, and abbreviation recovery. For noun phrase extraction, we employ the Berkeley Neural Parser (Kitaev and Klein, 2018, benepar) to obtain constituent trees for each sentence and recursively extract noun phrases. To handle coordinated expressions such as “broad and high nasal bridge”, which corresponds to both “broad nasal bridge” and “high nasal bridge” concepts, we introduce a coordinated phrase decomposition mechanism. Using spaCy’s dependency parser (Honnibal et al., 2020), we identify conjuncts for each coordinated token and then determine its longest constituent not overlapping with its conjuncts. Higher-level noun phrases covering multiple such constituents serve as templates to produce decomposed entities, where only one coordinated token’s constituent is used at once. This approach, compared with PhenoBERT which only handles A+B coordination, is capable of generalizing to more complex coordinated structures. Finally, we integrate the abbreviation detector from scispaCy (Neumann et al., 2019) to identify abbreviations and replace them by corresponding long forms, in order to reduce mismatch between abbreviations and unrelated identical concepts.
Definition 3 (Candidate concept retrieval).
Given an ontology O=(C,I)O{=}(C,I), an entity $e{=}T[i\mathord{:}j]$ from text $T$, and a maximum number of candidates $k$, the task is to find $f_{\text{CCR}}$ satisfying
$f_{\text{CCR}}(O,e,k)$ $=C_{\text{cand}}\subset C$ (3)
$\text{s.t.}\quad(|C_{\text{cand}}|\leq k,\;e\sim\exists c\in C_{\text{cand}}$ $)\vee(C_{\text{cand}}=\emptyset,\;e\not\sim\forall c\in C)$
which retrieves a small set of ontology concepts that potentially match each extracted entity.
To retrieve candidate concepts for each extracted entity $e$ from the input text, we adopt an embedding-based retrieval strategy with SapBERT, an embedding model trained to align UMLS concept names and synonyms. All ontology concepts and their synonyms are embedded in advance to construct a dense vector index, and then each extracted entity is embedded and compared against the index based on cosine similarity. If the highest similarity score exceeds a high-confidence threshold $\tau_{1}$, the entity is directly linked to the most similar concept. Otherwise, if the score falls within a relaxed interval $\tau_{2},\tau_{1})$, we retrieve the top-$k$ most similar concepts to form the candidate set $C_{\text{cand}}$ for downstream LLM-based entity linking. This hierarchical approach ensures high precision for unambiguous matches while preserving recall by delegating ambiguous cases to the more capable LLM.
Definition 4 (Entity linking).
Given an ontology O=(C,I)O{=}(C,I), an entity $e{=}T[i\mathord{:}j]$ from text $T$, and candidate concepts $C_{\text{cand}}$, the task is to find $f_{\text{EL}}$ satisfying
$f_{\text{EL}}(O,e,C_{\text{cand}})=c\in C_{\text{cand}}\sim e$ (4)
which selects the most semantically similar concept from the retrieved candidate concepts.
Inspired by the prompt-based method REAL, we perform entity linking by prompting an LLM. For each extracted entity $e$ and its retrieved candidate concept set $C_{\text{cand}}$, we construct a structured prompt that includes the entity name along with the list of candidate concepts. Each concept is represented by its ID, name, definition, synonyms, and cross-referenced UMLS synonyms. The output contains a concept ID or “None” with a confidence level, and only predictions with “HIGH” confidence are retained to reduce false positives. Specifically, the following prompt template is used.
System prompt:
As an expert clinician, your task is to accurately link the entity using the concepts listed below. Accuracy is paramount. If the entity does not precisely refer to any of the concepts listed below, please return “None”; otherwise, return the corresponding concept ID in the following format:
answer:<<concept ID or None>>
confidence:<<one of HIGH, LOW, MEDIUM>>
Below are the concepts:
{candidate concepts with ID, name, definition, synonyms, and UMLS synonyms.}
User prompt:
[Entity to link]
label: {entity string}
To facilitate the LLM-based entity linker to distinguish between highly similar concepts within the ontology, we introduce a self-supervised fine-tuning strategy. We first generate challenging training examples by leveraging the internal structure of the ontology and the retrieval capabilities of SapBERT. For each ontology concept, we perform synonym cleaning by lemmatizing all associated synonyms and merging those that become identical. For each unique lemmatized synonym $s$ belonging to a ground-truth concept $c_{gt}$, we use SapBERT to retrieve the top-$r$ non-repetitive candidate concepts. We then construct positive and negative training examples as follows: To create a positive candidate set $C_{\text{cand}}^{+}$, we check if $c_{gt}$ is present in the retrieved list. If it is, the list is preserved. If not, we replace the last concept in the list with $c_{gt}$ to ensure the ground truth is present among difficult distractors. Note that $s$ should be removed from the synonym list of $c_{gt}$ to avoid leakage. To create a negative candidate set $C_{\text{cand}}^{-}$, we ensure $c_{gt}$ is absent. If $c_{gt}$ is in the list, we remove it and append the (r+1)(r{+}1)-th retrieved concept to the end. If not, we retain the list as is. This procedure yields two candidate sets of length $r$ for each synonym, representing the most difficult cases for SapBERT to resolve. We transform these examples into training instances by applying the prompt template above with two modifications: (1) we remove the “confidence” and “UMLS synonyms” requirements from the system prompt to simplify the learning objective, and (2) we set the desired output to the ground-truth concept ID for positive examples and “None” for negative examples. By iterating through all concept synonyms, we generate a total of 42,110 training instances. We fine-tune the entity linker using QLoRA (Dettmers et al., 2023) for one epoch with $r=3$ candidates to achieve efficient adaptation, and name the resulting model $\text{AutoPCR}_{\text{FT}}$ from now on.
For post-processing, AutoPCR may produce overlapping entity spans linked to the same or different concepts. When linked to different concepts, all spans are retained. When linked to the same concept, the span with the highest similarity score is prioritized over those linked by the LLM. AutoPCR also provides an option to keep only the longest spans, similar to other neural methods, for scenarios where nested spans are not required.
4 Experiments
[…]
4.1 Benchmarks and baselines
To evaluate the performance of AutoPCR, we conduct experiments on four widely-used datasets: BIOC-GS, the development set from BioCreative VIII Track 3 (Weissenbacher et al., 2023), which consists of 382 clinical observation records from dysmorphology physical examinations, with an average length of 8.5 words. An example record is “ABDOMEN: Small umbilical hernia. Mild distention. Soft.” GSC-2024, a refined GSC+ (Lobo et al., 2017) dataset given by FastHPOCR, which comprises 228 PubMed abstracts, with an average length of 150 words. We use 22 of them for development, following PhenoTagger’s data split. ID-68, 68 real clinical notes from families with intellectual disabilities (Anazi et al., 2017), with an average length of 157 words. It is manually annotated by PhenoBERT in the same way as GSC+. NCBI (Doğan et al., 2014), consisting of 100 PubMed abstracts, with an average length of 205 words. BIOC-GS, GSC-2024, and ID-68 are grounded on HPO release 2024-02-08 under the root node “phenotypic abnormality” with 17,542 concepts. NCBI is grounded on MEDIC (Davis et al., 2025), a curated disease ontology integrating MeSH (Lipscomb, 2000) and OMIM (Amberger et al., 2019), under the root node “diseases” with 13,315 concepts. These datasets involve both free-form and standardized concept mentions, covering a wide range of use cases.
We follow the evaluation pipeline of PhenoTagger and PhenoTagger++, except that each predicted and gold entity is counted at most once to ensure one-to-one matching. Mention- and document-level precision (P%), recall (R%), and F1% are used as evaluation metrics. For mention-level evaluation, a predicted entity is considered correct only if there exists a ground-truth entity with the same linked concept and overlapping offsets. For document-level evaluation, linked concepts are treated as a set per document, and metrics are computed by comparing these sets, followed by micro-averaging over all documents. Average results over three runs are reported for all experiments.
We compare AutoPCR with competitive baselines from three categories: (1) dictionary-based methods, including NCBO, OBO, ClinPhen, and FastHPOCR; (2) neural methods, including NCR, PhenoTagger, PhenoBERT, and PhenoTagger++; and (3) prompt-based methods, including vanilla prompting (Groza et al., 2024a) and REAL. All baselines are tested with their recommended parameters. Prompt-based methods employ GPT-4o-mini as the LLM backend.
4.2 Implementation details
For entity extraction, we use Stanza packages “ner-i2b2” and “mimic”, spaCy model “en_core_web_trf”, benepar model “benepar_en3_large”, and scispaCy model “en_core_sci_scibert”. We recover abbreviations on ID-68, as they are counted separately. For candidate concept retrieval, we set $\tau_{1}{=}0.95{\in}\{0.9,0.925,0.95,0.975,1\}$ and $\tau_{2}{=}0.85{\in}\{0.8,0.825,0.85,0.875,0.9\}$ to distinguish high- and low-confidence matches. For similarity scores in $\tau_{2},\tau_{1})$, we retain up to $k{=}5{\in}\{3,5,7,9\}$ candidates. These hyperparameters are tuned on the GSC-2024 development set. For entity linking, we use Qwen3-Next-80B-A3B-Instruct as the frozen LLM backend and Qwen3-30B-A3B-Instruct-2507 as the base for fine-tuning (Yang et al., 2025). Temperature is fixed at zero for stability.
4.3 Main results
Table 1: Mention-level results on three datasets. AutoPCR is the highest in average F1 and F1 ranking.
Method BIOC-GS GSC-2024 ID-68 Avg. F1 Avg. Rk.
P R F1 (Rk.) P R F1 (Rk.) P R F1 (Rk.)
Dictionary- Based NCBO 82.20 45.63 58.69 (8) 95.71 51.57 67.03 (7) 86.27 65.23 74.29 (7) 66.67 7.33
OBO 71.94 42.67 53.57 (10) 85.89 52.66 65.29 (8) 81.41 60.79 69.61 (9) 62.82 9.00
ClinPhen 64.64 50.91 56.96 (9) 84.10 38.74 53.05 (10) 72.92 60.33 66.03 (10) 58.68 9.67
FastHPOCR 66.73 59.47 62.89 (5) 88.43 79.25 83.59 (1) 87.23 71.76 78.75 (3) 75.08 3.00
Neural NCR 64.04 57.50 60.59 (7) 68.54 74.88 71.57 (6) 78.64 78.18 78.41 (4) 70.19 5.67
PhenoTagger 71.12 65.73 68.32 (2) 86.16 78.07 81.92 (4) 84.40 72.58 78.04 (5) 76.09 3.67
PhenoBERT 74.77 53.71 62.51 (6) 85.90 74.88 80.01 (5) 93.36 78.76 85.44 (1) 75.99 4.00
PhenoTagger++ 69.16 63.92 66.44 (4) 87.71 78.22 82.69 (3) 79.90 72.81 76.19 (6) 75.11 4.33
Prompt- Based Vanilla 2.91 2.31 2.57 (11) 15.31 7.57 10.13 (11) 18.29 11.20 13.89 (11) 8.87 11.00
REAL 75.68 59.97 66.91 (3) 76.19 47.35 58.40 (9) 76.33 65.46 70.48 (8) 65.26 6.67
AutoPCR 84.25 67.27 74.81 (1) 87.13 79.19 82.97 (2) 83.68 76.59 79.98 (2) 79.25 1.67
$\text{AutoPCR}_{\text{FT}}$ 83.44 65.02 73.09 89.08 77.94 83.14 87.84 77.01 82.07 79.43
Table 2: Document-level results on three datasets. AutoPCR is the highest in average F1 and F1 ranking.
Method BIOC-GS GSC-2024 ID-68 Avg. F1 Avg. Rk.
P R F1 (Rk.) P R F1 (Rk.) P R F1 (Rk.)
Dictionary- Based NCBO 82.64 45.70 58.85 (8) 99.33 52.26 68.49 (7) 88.75 64.69 74.84 (7) 67.39 7.33
OBO 73.93 42.72 54.15 (10) 85.98 52.40 65.12 (8) 82.84 60.28 69.78 (9) 63.01 9.00
ClinPhen 64.78 51.16 57.17 (9) 86.20 44.99 59.12 (10) 74.96 61.92 67.82 (10) 61.37 9.67
FastHPOCR 70.92 59.77 64.87 (5) 95.15 77.61 85.49 (1) 87.75 71.38 78.72 (3) 76.36 3.00
Neural NCR 65.54 57.62 61.32 (7) 81.88 73.09 77.24 (6) 79.18 77.68 78.42 (4) 72.33 5.67
PhenoTagger 71.33 65.89 68.50 (2) 87.57 76.62 81.73 (4) 84.16 72.38 77.83 (5) 76.02 3.67
PhenoBERT 76.53 53.97 63.30 (6) 90.33 73.87 81.27 (5) 94.11 78.56 85.64 (1) 76.74 4.00
PhenoTagger++ 69.53 64.24 66.78 (4) 89.38 76.70 82.55 (3) 79.81 72.76 76.12 (6) 75.15 4.33
Prompt- Based Vanilla 4.42 3.48 3.89 (11) 18.25 9.75 12.71 (11) 20.04 11.98 15.00 (11) 10.53 11.00
REAL 76.47 60.27 67.41 (3) 80.21 54.66 65.02 (9) 76.64 66.20 71.04 (8) 67.82 6.67
AutoPCR 84.57 67.44 75.04 (1) 89.36 78.32 83.48 (2) 84.33 76.71 80.34 (2) 79.62 1.67
$\text{AutoPCR}_{\text{FT}}$ 83.94 65.18 73.38 91.62 76.69 83.50 88.77 77.05 82.49 79.79
Table 3: Mention-level recall on coordinated entities from the BIOC-GS, GSC-2024, and ID-68 datasets. AutoPCR achieves the highest recall.
Method Coordinated Entities
Dictionary- Based NCBO 5.98
OBO 26.63
ClinPhen 13.04
FastHPOCR 9.24
Neural NCR 15.76
PhenoTagger 19.57
PhenoBERT 21.74
PhenoTagger++ 25.00
Prompt- Based Vanilla 1.09
REAL 38.59
AutoPCR (w/o Coord.) 22.28
AutoPCR 45.11
The mention- and document-level results across three datasets grounded on HPO are shown in Tables 1 and 2, respectively. The two evaluation levels yield identical method rankings. AutoPCR ranks first on BIOC-GS and second on both GSC-2024 and ID-68, achieving the highest average F1 and the best overall rank. It also exhibits the most stable performance, demonstrating strong robustness across datasets. FastHPOCR achieves the strongest performance among dictionary-based methods, with the second-best average rank. Nevertheless, its performance drops notably on BIOC-GS, indicating reduced robustness on noisier clinical narratives. Traditional alternatives such as NCBO, OBO, and ClinPhen maintain high precision but suffer from low recall, resulting in overall lower F1. Neural methods PhenoTagger and PhenoBERT rank third and fourth overall. PhenoTagger performs better on BIOC-GS but worse on ID-68, whereas PhenoBERT shows the opposite trend. Both methods exhibit higher variability compared with AutoPCR. The earlier NCR is outperformed by its successor PhenoBERT, as expected, while PhenoTagger++, intended as a refinement of PhenoTagger, shows only a slight improvement on GSC-2024 but degraded performance on BIOC-GS and ID-68, leading to an overall decline and suggesting possible overfitting. Vanilla prompting performs poorly compared with all other baselines. In contrast, the RAG-based REAL performs strongly on BIOC-GS but ranks near the bottom on the others, highlighting the limitations of its generalist model design when applied to long, standardized text. These results comprehensively answer (RQ-1), demonstrating the superior and robust performance of AutoPCR across datasets. We further analyze these outcomes in detail, explaining how the design of AutoPCR contributes to its performance under different dataset characteristics.
On BIOC-GS consisting of noisy clinical notes, we observe a clear trend across method categories: prompt-based methods outperform neural methods, which in turn surpass dictionary-based ones, consistent with prior findings (Groza et al., 2024a; Qi et al., 2024; Shlyk et al., 2024). This pattern reflects their respective modeling capabilities. Prompt-based methods use LLMs with strong language understanding abilities for entity linking, enabling them to better handle free-form and lexically diverse clinical text. Neural methods, typically fine-tuned on structured, ontology-aligned corpora, struggle to generalize to such noisy input. Dictionary-based methods perform worst, as they lack semantic understanding and rely solely on exact or approximate string matching.
On GSC-2024 and ID-68, which feature more standardized text with less surface-level variation in entities, AutoPCR remains highly competitive, even though the advantage of language understanding becomes less pronounced. The top-performers on these two datasets, FastHPOCR and PhenoBERT, does not bring the advantage to the other; by contrast, AutoPCR consistently ranks behind them in the second place. This may be possibly because both methods partially annotated the datasets themselves, introducing inductive bias in their favor. Specifically, ID-68, as annotated by PhenoBERT, contains mentions that are not exactly matched by ontology concepts but instead linked to hypernyms (e.g., “thinning of the body of the corpus callosum” linked to HP:0002079 “Hypoplasia of the corpus callosum”). While this criterion may be helpful in certain use cases, it diverges from the standard task definition shared by our study and other competitive baselines. This inconsistency creates an ambiguous annotation standard: if every term covered by the ontology’s hierarchy is a potential candidate, the task essentially shifts into a classical hierarchical classification problem. Conversely, if only a subset is chosen, the labels become heavily influenced by the idiosyncratic styles of the annotators, making the problem more suitable for supervised learning rather than the zero-shot setting we address. This is also reflected in the experimental results, where PhenoBERT achieves an outlier precision on ID-68, hinting at a bias in the dataset’s construction. Nevertheless, AutoPCR still secures the second best, outperforming the next best baseline by approximately 3% in F1. Similarly, AutoPCR maintains a clear advantage over all other baselines except FastHPOCR on GSC-2024. This consistent lead demonstrates AutoPCR’s superior robustness, even when faced with deviating annotation criteria. Unlike on BIOC-GS, AutoPCR here significantly outperforms REAL, which can be credited to two novel designs: (1) a comprehensive entity extraction approach that improves recall on long, standardized text compared to LLM-based extraction and (2) a candidate retrieval module that effectively models domain-specific semantic similarity to filter noisy entities while prioritizing high-quality candidates.
The findings for AutoPCR are largely consistent with those of $\text{AutoPCR}_{\text{FT}}$, as they have a similar average F1 and share identical rankings if ranked separately. However, by fine-tuning on HPO with mostly standardized synonyms, $\text{AutoPCR}_{\text{FT}}$ trades off some of its ability to handle free-form text for enhanced precision in linking standardized entities. This shift is evidenced by a slight decrease in F1 on BIOC-GS, contrasted with a precision gain on GSC-2024 and ID-68. Specifically, on ID-68, both recall and precision improve simultaneously, leading to an F1 increase of over 2.0. These results validate our self-supervised fine-tuning strategy and demonstrate the effectiveness of ontology-specific adaptation when the input text style is aligned with the ontology’s.
To further evaluate each method’s ability to handle coordinated phrases, we curate 184 gold-standard entities from the BIOC-GS, GSC-2024, and ID-68 datasets that contain conjunctions such as “and”, “or”, “/”, “and/or”, or commas. These entities represent cases where a single phrase encodes multiple coordinated concepts, providing a focused benchmark for testing coordinated phrase decomposition. As shown in Table 3, AutoPCR achieves the highest recall, while its variant without coordinated phrase decomposition performs comparably to other baselines. This confirms the effectiveness of our decomposition mechanism in correctly decomposing coordinated entities into separate concepts. REAL ranks second, as the LLM-based entity extraction can inherently split coordinated mentions, but it still falls short of AutoPCR and generates substantially more output tokens, leading to higher computational and monetary costs. Among neural methods, PhenoBERT, despite claiming to handle coordination like “A+B”, does not show noticeable improvement over its peers. Within dictionary-based methods, OBO stands out, which is expected since it heuristically concatenates noncontiguous subsequences into entities, capturing simple coordinated cases. To summarize, a representative example may be the phrase “hearing loss of conductive, sensorineural, or mixed type”, which requires decomposition into three distinct concepts: “hearing loss of {conductive, sensorineural, mixed} type”. Only AutoPCR can achieve this.
Ablation study
Table 4: Ablation study of AutoPCR on syntactic entity extraction, HPO/UMLS knowledge, and entity linking. Removing any component overall degrades performance. mF1: mention-level F1; dF1: document-level F1.
Syntactic Entities HPO Knowledge UMLS Knowledge Entity Linking BIOC-GS GSC-2024 ID-68 Avg. mF1 Avg. dF1
mF1 dF1 mF1 dF1 mF1 dF1
Variant 1 $\checkmark$ $\checkmark$ $\checkmark$ 67.58 67.84 82.18 82.94 79.38 79.60 76.38 76.79
Variant 2 $\checkmark$ $\checkmark$ $\checkmark$ 74.10 74.33 82.61 83.22 78.77 78.99 78.49 78.84
Variant 3 $\checkmark$ $\checkmark$ $\checkmark$ 16.83 17.13 18.71 21.87 20.54 25.09 18.69 21.36
Variant 4 $\checkmark$ $\checkmark$ $\checkmark$ 69.90 70.19 81.43 81.37 80.80 80.93 77.38 77.49
AutoPCR $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 74.81 75.04 82.97 83.48 79.98 80.34 79.25 79.62
To assess the contribution of each module in AutoPCR and answer (RQ-2), we conduct ablation experiments shown in Table 4. Specifically, we evaluate the following four variants of AutoPCR. Variant 1 removes syntactic entity extraction and relies solely on bioNER for entity extraction. It assesses whether syntactic entities contribute beyond direct BioNER segmentation. Variant 2 disables HPO knowledge and retrains SapBERT (initialized from PubMedBERT) on UMLS excluding all HPO-related concepts. It simulates a scenario where the target ontology is newly introduced and unseen during model training. Variant 3 examines the necessity of domain-specific alignment in SapBERT by replacing it with PubMedBERT for embedding, without any fine-tuning on UMLS. Variant 4 evaluates the importance of entity linking by entirely removing the linking module and rely solely on candidate retrieval for concept prediction. It makes prediction only if the top candidate concept exceeds the similarity threshold $\tau_{1}$.
The full AutoPCR model overall outperforms all ablated variants across datasets and metrics, confirming the importance of each component. Variant 1 shows a clear performance decline on BIOC-GS, confirming that BioNER alone is insufficient for handling free-form text and that syntactic entity extraction is essential for improving recall in such cases. Variant 3 exhibits the most severe performance drop, demonstrating that domain-specific alignment in embedding-based retrieval is critical for accurate concept grounding. Variant 4 performs reasonably well on standardized abstracts such as GSC-2024 and ID-68, suggesting that surface-form matching alone may suffice for low-ambiguity datasets. However, it underperforms markedly on BIOC-GS, which contains noisier and more ambiguous clinical narratives, underscoring the value of LLM-based reasoning for fine-grained semantic disambiguation. Notably, Variant 2 achieves the strongest results among all variants and remains comparable to the full AutoPCR model even without HPO-specific knowledge, indicating that AutoPCR can perform inductive inference on entirely new or rapidly evolving ontologies, such as HPO.
Generalizability
Table 5: Performance of selected methods on NCBI. AutoPCR obtains the highest accuracy with the lowest deployment time.
Method NCBI
mF1 dF1 Deployment Time (min)
FastHPOCR 49.54 57.11 4.2
NCR 34.24 31.80 1680
PhenoTagger 67.48 68.83 906
AutoPCR 73.15 71.93 2.6
$\text{AutoPCR}_{\text{FT}}$ 72.76 70.43 2.6
To evaluate the generalizability and deployment efficiency of AutoPCR, we conduct an additional experiment on the NCBI dataset using the MEDIC ontology, as shown in Table 5. We define deployment time as the total time needed to prepare a method for inference on a new ontology, including ontology-specific index construction or model retraining. We compare AutoPCR with three strong and representative baselines: FastHPOCR, NCR, and PhenoTagger. PhenoTagger++ and PhenoBERT are excluded due to the lack of publicly available pipelines for adapting to a new ontology.
[…]
Hyperparameter sensitivity
Table 6: Performance of AutoPCR with various open-source and proprietary LLM backends.
LLM Backend BIOC-GS GSC-2024 ID-68 NCBI Avg. mF1 Avg. dF1
mF1 dF1 mF1 dF1 mF1 dF1 mF1 dF1
Open- Source Llama-3.1-8B-Instruct 70.86±0.32 71.06±0.32 79.93±0.08 79.83±0.10 78.50±0.13 78.74±0.12 69.75±0.04 65.62±0.04 74.76±0.14 73.81±0.15
Llama-3.3-70B-Instruct 72.63±0.19 72.85±0.20 82.33±0.04 82.72±0.05 79.20±0.11 79.56±0.13 71.00±0.02 68.07±0.14 76.29±0.09 75.80±0.13
Llama-4-Scout-17B-16E-Instruct 72.41±0.25 72.70±0.25 82.14±0.22 82.25±0.31 80.90±0.12 81.16±0.11 72.93±0.21 70.74±0.22 77.09±0.20 76.71±0.23
Qwen3-30B-A3B-Instruct-2507 72.60±0.03 72.81±0.03 81.71±0.01 82.03±0.00 78.73±0.00 79.07±0.00 71.36±0.02 68.84±0.00 76.10±0.01 75.69±0.01
Qwen3-Next-80B-A3B-Instruct 74.81±0.13 75.04±0.13 82.97±0.09 83.48±0.11 79.98±0.11 80.34±0.11 73.15±0.18 71.93±0.34 77.73±0.13 77.70±0.17
DeepSeek-V3.1 73.66±0.21 73.88±0.21 83.15±0.03 83.80±0.09 80.14±0.07 80.42±0.07 72.47±0.02 71.21±0.08 77.35±0.08 77.33±0.11
Kimi-K2-Instruct-0905 72.77±0.12 72.98±0.12 83.00±0.25 83.70±0.31 79.33±0.14 79.74±0.17 71.46±0.37 68.70±1.23 76.64±0.22 76.28±0.46
Proprietary GPT-4o-mini 73.95±0.11 74.18±0.11 82.71±0.09 83.04±0.10 79.36±0.10 79.63±0.11 71.54±0.05 69.05±0.07 76.89±0.09 76.47±0.10
GPT-4.1-nano 72.55±0.24 72.76±0.24 81.92±0.06 81.93±0.10 79.27±0.12 79.55±0.13 71.06±0.10 68.54±0.15 76.20±0.13 75.69±0.16
GPT-4.1-mini 73.66±0.07 73.88±0.07 82.67±0.03 82.90±0.02 79.18±0.09 79.55±0.12 71.79±0.09 68.80±0.09 76.83±0.07 76.28±0.07
GPT-4.1 74.41±0.19 74.63±0.19 83.12±0.07 83.64±0.11 79.46±0.05 79.84±0.05 72.28±0.06 69.91±0.17 77.32±0.09 77.00±0.13
Figure 2: Sensitivity analysis of AutoPCR w.r.t. $\tau_{1}$ shows robustness.
Figure 3: Sensitivity analysis of AutoPCR w.r.t. $\tau_{2}$ and $k$ shows robustness.
Lastly, we examine the sensitivity of key hyperparameters in AutoPCR, including the high-confidence threshold $\tau_{1}$, the low-confidence threshold $\tau_{2}$, the number of retrieved candidates $k$, and the choice of LLM backends, on the GSC-2024 dataset.
Figures 2 and 3 show the effects of varying $\tau_{1}$, $\tau_{2}$, and $k$. For the high-confidence threshold $\tau_{1}$, performance improves slightly as $\tau_{1}$ increases from 0.9 to 0.95, likely because more false positives are filtered by the LLM-based entity linking stage. When $\tau_{1}$ exceeds 0.95, performance begins to decline and drops sharply at $\tau_{1}{=}1.0$, where all entities above the low-confidence threshold are forced into the linking stage. This indicates that the LLM can introduce false negatives when surface-form matches are already sufficient, underscoring the need to retain a high-confidence threshold for direct predictions. Overall, AutoPCR is robust to $\tau_{1}$ across the range of 0.9–0.975, with the default $\tau_{1}{=}0.95$ achieving the best F1. When varying $\tau_{2}$ and $k$, which jointly determine the number and quality of candidate concepts for entity linking, performance is largely stable for $\tau_{2}{\in}[0.8,0.875]$, peaking at the default $\tau_{2}{=}0.85$. Smaller $\tau_{2}$ introduces excessive noise by retrieving too many irrelevant candidates, whereas larger $\tau_{2}$ risks missing valid ones, both leading to a degradation in F1. On the other hand, increasing $k$ from 3 to 7 yields marginal performance gains at the expense of increased computational overhead, representing a typical trade-off between accuracy and efficiency.
Table 6 summarizes the performance of AutoPCR across eleven representative non-reasoning LLM backends (Grattafiori et al., 2024; Meta, 2025; Liu et al., 2024; Team et al., 2025) under three random seeds. We have the following findings: (1) The choice of LLM family has only a modest effect on performance. Flagship models from different families, such as Llama-4-Scout, Qwen3-Next-80B, DeepSeek-V3.1, and GPT-4.1, exhibit a performance delta of less than 1.0 F1, demonstrating the robustness of AutoPCR across various LLM backends. (2) Random seeds have negligible impact. The variance in F1 remains below 0.2 for most backends except Kimi-K2 and Llama-4-Scout, confirming the stability of AutoPCR. (3) Performance generally scales with model size. Within the Llama family, F1 scores rise steadily from the 3B to the 70B and 109B (Scout) models. Similar upward trends are observed across the GPT family (from nano to mini to GPT-4.1) and the Qwen series (from 30B to 80B). The large-scale DeepSeek-V3.1 (685B) also performs strongly, ranking near the top. These results highlight the scalability of AutoPCR and its ability to benefit from increasingly capable LLMs. (4) For practical deployment, the open-source Qwen3-Next-80B offers the best overall performance with reasonable inference efficiency, while for proprietary settings, smaller GPT-mini models provide an excellent trade-off between accuracy and cost.
5 Conclusion
In this work, we have developed AutoPCR, an automated phenotype CR method by prompting. Our experiments show that AutoPCR not only achieves superior accuracy and robustness on HPO, but also exhibits strong inductive capability and generalizability to an unseen ontology, hence the name “automated”.
6 Conflicts of interest
The authors declare that they have no competing interests.
7 Funding
This work is supported by NIH #U24DK138515 and #OT2OD038003.
8 Data availability
Datasets and ontologies are sourced from Weissenbacher et al. (2023, BIOC-GS), Groza et al. (2024b, GSC-2024), Feng et al. (2022, ID-68), Luo et al. (2021, NCBI), Talapova et al. (2023, HPO), and Davis et al. (2025, MEDIC). Access links are available in our code repository.
9 Author contributions statement
Y.T., Y.H., and J.L. conceived the methodology. Y.T. and Y.W. conducted the experiments. Y.T., Y.H., Y.W., and X.L. analyzed the results. Y.T. and J.L. wrote and reviewed the manuscript.